# Limits of Absoluteness of Observed Events in Timelike Scenarios: A No-Go Theorem

### v1

@@ III.3 Proof of the Causal Friendliness No-Go Theorem
Consequence of ATS: From Bayes’ rule we can write,
@@
If AOE, ATS and NRC hold, then Charlie and Debbie’s joint outcome statistics are independent of Alice and Bob’s choice of measurements $x$ and $y$, i.e.,
@@
From the assumption of ATS we have,
@@
Finally, combining Eqs. (17) and 18 with Eq. (16) gives,
@@
If AOE, ATS and NRC hold, then Alice’s(Bob’s) outcome aa($b$) is effectively screened off by the Pseudo Events in its causal past; that is, it cannot depend on Bob’s (Alice’s) choice of measurement, and hence,
@@
However, due to NRC we can always write $p(a|c,x,d)=p(a|c,x)$.
Similarly, for the time-reversed scenario, using the ATS assumption (for notational simplicity $^{\prime}\rightarrow^{\prime}$ in the subscript is omitted), we can write,
@@
We have now established two important results in the form of Lemmas 1 and 2, which will be used to prove the main result. At this stage, we insert the SPE assumption (i.e., p(b|c,x,a,d,y)=p(b|c,x,d,y)b,c,x,d,yp(b|c,x,a,d,y)=p(b|c,x,d,y)\ \ $\forall$ b,c,x,d,y) into the causal factorisation (Eq. (12)), then inserting Lemmas 1 and 2, and NRC assumption to the modified expression yields
@@
Therefore, no theory satisfying AOE and Causal Agency can reproduce all quantum Causal Friendliness correlations. ∎
IV Role of Pseudo Events in the Causal Friendliness Paradox
From the analysis in the previous section, a natural question arises: does the violation of Eq. (35) imply that the observed events are not absolute (even the Pseudo Events), but instead subjective to the observers? To address this, we first examine to what extent the AOE assumption is truly necessary, in addition to ATS, NRC, and SPE, for establishing the Causal Friendliness no-go theorem. In this context, we attempt to supersede the AOE assumption by drawing an analogy with the replacement of the “wavefunctions are $\psi$-ontic” assumption used in Price’s no-go result [16], by the $\lambda$-mediation assumption (discussed later in this section), which Leifer and Pusey used to establish a no-go theorem [17] about time symmetry and retrocausal ontology. Price’s no-go theorem [16] suggests that any time-symmetric ontology of quantum theory must entail retrocausality. His argument proceeds by showing that the following three assumptions, when taken together, contradict quantum theory:
@@ IV Role of Pseudo Events in the Causal Friendliness Paradox
    Note this is defined analogously to condition NRC given earlier in this paper, but applied to hidden variable $\lambda$ rather than Pseudo Events $c$ and $d$;
@@
    $p_{\rightarrow}(\lambda,a,b\mid x,y)\;=\;p_{\leftarrow}(b,a,\lambda\mid y,x).$ (37)
    This is the the ontological counterpart of the ATS assumption, that is applied to probabilities associated with $\lambda$ instead of the Pseudo Events;
@@ IV.1 Replacing $\psi$-ontic reproducibility with $\lambda$-mediation
We here ask whether we can relax the AOE assumption to a weaker assumption in the same way as Leifer and Pusey replace Price’s $\psi$-ontic assumption with $\lambda$-mediation [17]. Given the AOE assumption seems similar to the $\psi$-ontic assumption (albeit imposed on the operational and observational level instead of the ontological/hidden-variable level), we argue that the answer is “yes”. To show this is the case, we first examine formally the relation between the AOE assumption, and the $\psi$-ontic criterion for ontological models. We then show that the AOE assumption can be relaxed, to an operational equivalent of $\lambda$-mediation, which is weaker than AOE, while still allowing us to form Local Friendliness-style inequalities which quantum theory violates.
IV.2 Can we weaken the AOE assumption?
The logical structure underlying our weakening of the AOE assumption closely follows Leifer and Pusey’s weakening of Price’s $\psi$-ontic assumption to $\lambda$-mediation [17] - both replacing a strong absoluteness assumption ($\psi$-ontic reproducibility or AOE) with a weaker one (ontic or operational mediation). The key similarity between the arguments in [17] and ours lies in the fact that the independence of the mediator variables from the choice of measurement settings are not assumed but rather derived from other underlying assumptions.
Theorem 2 (Operational Pseudo Event Mediation (OPEM)).
Let the Causal-Friendliness scenario be as in Sec. III. We replace AOE (existence of the full joint $p(c,a,d,b|x,y)$) with the following weaker, purely operational assumptions:
@@ IV.2 Can we weaken the AOE assumption?
    Existence of marginals (EOM): For all settings $(x,y)$ there exist well-defined operational marginals $p(c,d|x,y)$ and conditional response functions $p(a|x,c)$ and $p(b|y,c,d)$ describing how the outcomes of Truly-Observed Events $a$ and $b$ depend on Pseudo Events and choice of measurement settings. (Note we make no claim about existence of the full four-variable joint probability for each run.)
@@
    Operational mediation (OM): Pseudo Events $c,d$ acts as the mediator between future and past events such that once $c,d$ are specified, $a$ provides no further information on $b$ beyond that already provided by $c,d$ and $y$, i.e.,
@@
Combining these with the ATS and NRC assumptions (as in Sec. III), then the operational probabilities admit the factorisation
$p(a,b|x,y)\;=\;\sum_{c,d}p(c,d)\,p(a|x,c)\,p(b\mid y,c,d),$ (40)
and hence the Causal-Friendliness bound $|S|\leq 2$ follows.
@@
In order to prove the the factorisation of operational probability $p(a,b|x,y)$ as given in Eq. (40), we recall the operational ingredients posited by OPEM - specifically,
@@
Crucially, we do not assume the prior existence of a Kolmogorovian joint probability measure $p(a,b,c,d\mid x,y)$ defined by all the four events. Instead construct the probability from the operational marginals that are empirically defined. To this end we define such a operationally constructed distribution for operational joint probability-like object as
$q(a,b,c,d|x,y)\;:=\;p(c,d|x,y)\;p(a\mid x,c)\;p(b|y,c,d).$ (42)
The right-hand side of Eq. (42) uses only the operational objects that OPEM assumes exist. Therefore, $q$ should be regarded as a constructed function of these empirical distributions, rather than as a postulate implying the prior existence of a full joint distribution. However, we expect that for every fixed $(x,y)$, $q(\cdot,\cdot,\cdot,\cdot|x,y)$ is a non-negative function which sums up to unity as each factor constitutes a conditional probability:
@@
Hence $q(\cdot,\cdot,\cdot,\cdot\mid x,y)$ is a valid probability distribution. Defining $q$ in Eq. (42) is not the same as assuming the existence of the previously given Kolmogorovian joint probability. Rather, $q$ is built from marginals and conditionals which are assumed to be empirically defined entities; this is a strictly weaker commitment than postulating a global joint probability a priori.
@@
By construction $\widetilde{p}(a,b|x,y)$ is the operational prediction obtained by combining the assumed marginals in OPEM. To prove that $\widetilde{p}(a,b|x,y)$ is same as $p(a,b|x,y)$ given in Eq. (40) we have to prove the independence of $c,d$ from $x,y$. However, we cannot establish this result in the same manner as Lemma 1, since the proof of Lemma 1 relies on the assumption of AOE, which we do not use here. Using instead the operational mediation assumption of OPEM, we write the constructed distribution as in Eq. (42).
Further, ATS implies that the forward and time-reversed operational constructions coincide under the exchange $(c,a,x)\!\leftrightarrow\!(d,b,y)$, meaning we can write,
@@
where the quantities on the right-hand side are again operational marginals defined for the time-reversed experiment. Note that, every object that appears in the algebra above (namely, the factors in $q$ and the equalities given by ATS) is either an operational marginal (e.g. $(p(c,d|x,y)$) or an operational conditional (e.g. $p(a|x,c)$, $p(b|y,c,d)$), all assumed to exist by OPEM. We never invoked any property that would require these to be derived from the same pre-existing global Kolmogorovian probability distribution; rather we constructed $q$ from the operational probabilities. Equating Eqs. (42) and (45) gives
@@
Summing over $a$ and $b$ we get from Eq. (46) $p(c,d|x,y)=p(c,d|y,x)$. Finally, applying NRC in the forward and time reversed scenario, we get $p(c,d|x,y)=p(c,d|x)$ and $p(c,d|y,x)=p(c,d|y)$. This simply implies that they both are equal to $p(c,d)$, completing the proof. ∎
@@
It is important to note that the assumption OM is stronger than the SPE assumption used in the proof of the Causal-Friendliness no-go theorem. In other words, SPE can be viewed as a specific instance of Operational mediation, where the dependence of the event $b$ on the past input choice $x$ is not relaxed. In the proof of Causal-Friendliness no-go theorem, however, the independence of $b$ from $x$ is derived from the joint assumptions of AOE, ATS, and NRC. In contrast, in the present analysis, we are not permitted to assume AOE at the outset, and consequently, the independence of $b$ from $x$ is incorporated directly into the relatively stronger assumption of Operational mediation itself.
@@
Another important point to note here is that, although EOM by itself does not require the existing marginals to originate from a single global joint probability distribution, imposing the requirement that those existing marginals given by EOM originate from a single global joint probability distribution leads to AOE. This additional requirement can be summarised as follows:
Definition 2 (Existence of Joint Probability Distributions (EJPD)).
A theory is said to satisfy the Existence of Joint Probability Distributions if all marginals corresponding to the observed events can be recovered from a single global joint probability distribution $p(a,b,c,d|x,y)$.
Then by considering the above definition we have,
$\textbf{EOM}+\textbf{EJPD}\implies\textbf{AOE}$ (47)

### v2

@@ III.3 Proof of the Causal Time-Symmetric Friendliness No-Go Theorem
Consequence of Axiological Time Symmetry: From Bayes’ rule we can write,
@@
If Absoluteness of Observed Events, Axiological Time Symmetry and No Retrocausality hold, then Charlie and Debbie’s joint outcome statistics are independent of Alice and Bob’s choice of measurements $x$ and $y$, i.e.,
@@
From the assumption of Axiological Time Symmetry we have,
@@
Finally, combining Eqs. (19) and 20 with Eq. (18) gives,
@@
If Absoluteness of Observed Events, Axiological Time Symmetry and No Retrocausality hold, then Alice’s(Bob’s) outcome aa($b$) is effectively screened off by the Pseudo Events in its causal past; that is, it cannot depend on Bob’s (Alice’s) choice of measurement, and hence,
@@
However, due to No Retrocausality we can always write $p(a|c,x,d)=p(a|c,x)$.
Similarly, for the time-reversed scenario, using the Axiological Time Symmetry assumption (for notational simplicity $^{\prime}\rightarrow^{\prime}$ in the subscript is omitted), we can write,
@@
We have now established two important results in the form of Lemmas 1 and 2, which will be used to prove the main result. At this stage, we insert the Screening via Pseudo Events assumption (i.e., p(b|c,x,a,d,y)=p(b|c,x,d,y)b,c,x,d,yp(b|c,x,a,d,y)=p(b|c,x,d,y)\ \ $\forall$ b,c,x,d,y) into the causal factorisation (Eq. (14)), then inserting Lemmas 1 and 2, and No Retrocausality assumption to the modified expression yields
@@
Therefore, no theory satisfying Absoluteness of Observed Events and Causal Agency can reproduce all quantum Causal Time-Symmetric Friendliness correlations. ∎
III.4 Relation to the causal Markov condition
It is useful to clarify how the assumptions employed in this work relate to the standard causal Markov condition used in directed acyclic graph-based causal models [45]. For the forward time ordering of the Causal Time-Symmetric Friendliness scenario, a natural causal structure would suggest the Markov factorisation
$\begin{split}p&(c,a,d,b|x,y)=\\ &p(c)\,p(a|c,x)\,p(d|c,a)\,p(b|c,a,d,y),\end{split}$ (38)
from which all conditional independence relations would normally follow. In this work, however, we do not assume the causal Markov condition. Instead, we impose a restricted set of operationally-motivated conditional independence constraints, encoded in the assumptions of No Retrocausality and Screening via Pseudo Events. As a consequence, some of the conditional independences implied by the causal Markov condition are retained, while others are do not appear in our factorisation. Specifically, No Retrocausality enforces that future measurement choices influence neither past Pseudo Events nor past Truly-Observed Events. This retains Markov-type independences such as the independence of $c$ from $x,y$, and the independence of $a$ from $y$ once its past variables are fixed. Screening via Pseudo Events further ensures that future Truly-Observed Events are screened off from earlier events by appropriate Pseudo Events, thereby retaining the Markov style screening of $b$ by $c,d,y$.
Crucially, we do not require that conditioning on immediate causal parents always removes all further dependencies. In particular, we allow the Pseudo Event $d$ to depend on either Alice’s measurement choice nor her outcome, and we do not impose the minimal parent screening characteristic of the causal Markov condition. These dropped independences are essential for accommodating Pseudo Events, and for distinguishing our framework from standard hidden variable models or directed acyclic graph-based causal models. The relation between the conditional independences implied by the causal Markov condition and those assumed in this work is summarised in Table 1. Overall, our framework abandons the global causal Markov assumption and replaces it with a controlled set of conditional independence relations justified directly by No Retrocausality and Screening via Pseudo Events, which are sufficient for deriving the Causal Time-symmetric Friendliness inequality.
Conditional independence Causal Markov    Status here
$p(c|x,y)=p(c)$    Implied    Retained by No Retrocausality
$p(a|c,x,d,y)=p(a|c,x,d)$    Implied    Retained by Lemma 2
$p(b|a,c,d,x,y)=p(b|c,d,y)$    Not implied    Retained by Screening via Pseudo Events and Lemma 2
$p(d|c,a,x)=p(d|c,a)$    Implied    Does not appear
$p(b|c,a,d,y,x)=p(b|d,y)$    Implied    Does not appear
Table 1: Comparison between conditional independences implied by the causal Markov condition and those employed in this work.
IV Role of Pseudo Events in the Causal Time-Symmetric Friendliness Paradox
From the analysis in the previous section, a natural question arises: does the violation of Eq. (37) imply that the observed events are not absolute (even the Pseudo Events), but instead subjective to the observers? To address this, we first examine to what extent the Absoluteness of Observed Events assumption is truly necessary, in addition to Axiological Time Symmetry, No Retrocausality, and Screening via Pseudo Events, for establishing the Causal Time-Symmetric Friendliness no-go theorem. In this context, we attempt to supersede the Absoluteness of Observed Events assumption by drawing an analogy with the replacement of the “wavefunctions are $\psi$-ontic” assumption used in Price’s no-go result [48], by the $\lambda$-mediation assumption (discussed later in this section), which Leifer and Pusey used to establish a no-go theorem [39] about time symmetry and retrocausal ontology. Price’s no-go theorem [48] suggests that any time-symmetric ontology of quantum theory must entail retrocausality. His argument proceeds by showing that the following three assumptions, when taken together, contradict quantum theory:
@@ IV Role of Pseudo Events in the Causal Time-Symmetric Friendliness Paradox
    Note this is defined analogously to condition No Retrocausality given earlier in this paper, but applied to hidden variable $\lambda$ rather than Pseudo Events $c$ and $d$;
@@
    $p_{\rightarrow}(\lambda,a,b|x,y)\;=\;p_{\leftarrow}(b,a,\lambda|y,x).$ (40)
    This is the the ontological counterpart of the Axiological Time Symmetry assumption, that is applied to probabilities associated with $\lambda$ instead of the Pseudo Events;
@@ IV.1 Replacing $\psi$-ontic reproducibility with $\lambda$-mediation
We here ask whether we can relax the Absoluteness of Observed Events assumption to a weaker assumption in the same way as Leifer and Pusey replace Price’s $\psi$-ontic assumption with $\lambda$-mediation [39]. Given the Absoluteness of Observed Events assumption seems similar to the $\psi$-ontic assumption (albeit imposed on the operational and observational level instead of the ontological/hidden-variable level  [^1] Quantum states encode probabilities for all possible events, typically corresponding to measurement outcomes. In a $\psi$-ontic model, distinct quantum states correspond to ontic states with disjoint support in the underlying ontic state space. Analogously, the assumption of Absoluteness of Observed Events asserts that distinct observed events correspond to disjoint support in the relevant probability space. Thus, although the assumptions are structurally and physically different, Absoluteness of Observed Events plays a role analogous to $\psi$-onticness when one is concerned with the definiteness and global consistency of an underlying description.), we argue that the answer is “yes”. To show this is the case, we first examine formally the relation between the Absoluteness of Observed Events assumption, and the $\psi$-ontic criterion for ontological models. We then show that the Absoluteness of Observed Events assumption can be relaxed, to an operational equivalent of $\lambda$-mediation, which is weaker than AOE, while still allowing us to form Local Friendliness-style inequalities which quantum theory violates.
IV.2 Can we weaken the Absoluteness of Observed Events assumption?
The logical structure underlying our weakening of the Absoluteness of Observed Events assumption closely follows Leifer and Pusey’s weakening of Price’s $\psi$-ontic assumption to $\lambda$-mediation [39] - both replacing a strong absoluteness assumption ($\psi$-ontic reproducibility or Absoluteness of Observed Events) with a weaker one (ontic or operational mediation). The key similarity between the arguments in [39] and ours lies in the fact that the independence of the mediator variables from the choice of measurement settings are not assumed but rather derived from other underlying assumptions.
Let the Causal Time-Symmetric Friendliness scenario be as in Sec. III. We replace Absoluteness of Observed Events (existence of the full joint $p(c,a,d,b|x,y)$) with the following weaker, purely operational assumption:
Definition 5 (Operational Pseudo Event Mediation ).
A theory admits Operational Pseudo Event Mediation if all of its statistics corresponding any arbitrary time-ordered prepare-measure scenario satisfies the following two assumptions.
@@ IV.2 Can we weaken the Absoluteness of Observed Events assumption?
    Existence of marginals: For all measurements there exist well-defined operational marginals for the joint probabilities of the Pseudo Events and conditional response functions for the Truly-Observed Events that are independent of all the measurements settings excepts its immediate cause, describing how the outcomes of Truly-Observed Events depend on Pseudo Events and choice of measurements (The assumption does not require that these response functions and the joint probability distribution for the Pseudo Events arise from a single global probability distribution.)
    For Causal Time-Symmetric Friendliness scenario this implies that for all settings $(x,y)$ there exist operational marginals $p(c,d|x,y)$ and conditional response functions $p(a|x,c)$ and $p(b|y,c,d)$.
@@
    Operational mediation: All the Pseudo Events act as mediators between past and future events, such that once they are specified, one Truly-Observed Event provides no information about a future Truly-Observed Event beyond that already provided by the Pseudo Events in its past and the measurement that is its immediate cause.
    The specific version of this assumption, tailored to the present scenario, can be written technically as follows:
@@
The above definition can be summarised as the following implication,
$\textbf{Existence of marginals}+\textbf{Operational mediation}$
$\ \implies\textbf{Operational Pseudo Event Mediation}$ (43)
With these assumptions in place, we establish the following result.
Theorem 2.
Any theory that admits Operational Pseudo Event Mediation, together with the Axiological Time Symmetry and No Retrocausality assumptions (as defined in Sec. III), yields the following factorisation of the operational probabilities
$p(a,b|x,y)\;=\;\sum_{c,d}p(c,d)\,p(a|x,c)\,p(b|y,c,d),$ (44)
and hence the Causal Time-Symmetric Friendliness bound $|S|\leq 2$ follows.
@@
In order to prove that the factorisation of operational probability $p(a,b|x,y)$ as given in Eq. (44), we recall the operational ingredients posited by Operational Pseudo Event Mediation - specifically,
@@
Crucially, we do not assume the prior existence of a Kolmogorovian joint probability measure $p(a,b,c,d|x,y)$ defined by all the four events. Instead construct the probability from the operational marginals that are empirically defined. To this end we define such a operationally constructed distribution for operational joint probability-like object as
q(a,b,c,d|x,y):=p(c,d|x,y)p(a|x,c)p(b|y,c,d).q(a,b,c,d|x,y)\;:=\;p(c,d|x,y)\;p(a|x,c)\;p(b|y,c,d). (46)
The right-hand side of Eq. (46) uses only the operational objects that Operational Pseudo Event Mediation assumes exist. Therefore, $q$ should be regarded as a constructed function of these empirical distributions, rather than as a postulate implying the prior existence of a full joint distribution. However, we expect that for every fixed $(x,y)$, $q(\cdot,\cdot,\cdot,\cdot|x,y)$ is a non-negative function which sums up to unity as each factor constitutes a conditional probability:
@@
Hence $q(\cdot,\cdot,\cdot,\cdot|x,y)$ is a valid probability distribution. Defining $q$ in Eq. (46) is not the same as assuming the existence of the previously given Kolmogorovian joint probability. Rather, $q$ is built from marginals and conditionals which are assumed to be empirically defined entities; this is a strictly weaker commitment than postulating a global joint probability a priori.
@@
By construction $\widetilde{p}(a,b|x,y)$ is the operational prediction obtained by combining the assumed marginals in Operational Pseudo Event Mediation. To prove that $\widetilde{p}(a,b|x,y)$ is same as $p(a,b|x,y)$ given in Eq. (44) we have to prove the independence of $c,d$ from $x,y$. However, we cannot establish this result in the same manner as Lemma 1, since the proof of Lemma 1 relies on the assumption of Absoluteness of Observed Events, which we do not use here. Using instead the operational mediation assumption of Operational Pseudo Event Mediation, we write the constructed distribution as in Eq. (46).
Further, Axiological Time Symmetry implies that the forward and time-reversed operational constructions coincide under the exchange $(c,a,x)\!\leftrightarrow\!(d,b,y)$, meaning we can write,
@@
where the quantities on the right-hand side are again operational marginals defined for the time-reversed experiment. Note that, every object that appears in the algebra above (namely, the factors in $q$ and the equalities given by Axiological Time Symmetry) is either an operational marginal (e.g. $(p(c,d|x,y)$) or an operational conditional (e.g. $p(a|x,c)$, $p(b|y,c,d)$), all assumed to exist by Operational Pseudo Event Mediation. We never invoked any property that would require these to be derived from the same pre-existing global Kolmogorovian probability distribution; rather we constructed $q$ from the operational probabilities. Equating Eqs. (46) and (49) gives
@@
Summing over $a$ and $b$ we get from Eq. (50) $p(c,d|x,y)=p(c,d|y,x)$. Finally, applying No Retrocausality in the forward and time reversed scenario, we get $p(c,d|x,y)=p(c,d|x)$ and $p(c,d|y,x)=p(c,d|y)$. This simply implies that they both are equal to $p(c,d)$, completing the proof. ∎
We emphasize that the motivation for introducing Existence of Marginals andOperational Mediation is not merely to weaken assumptions mathematically, but to identify the minimal operational structure required to derive the Causal Time-Symmetric Friendliness inequalities. To illustrate the distinction between Absoluteness of Observed Events and the weaker assumptions of Existence of Marginals and Existence of Marginals, consider a prepare-measure scenario in which certain intermediate events may be unitarily erased by a later intervention. In such a case, the operationally accessible statistics consist of well defined marginals, such as $p(c,d|x,y)$, together with response functions $p(a|x,c)$ and $p(b|y,c,d)$ associated with Truly-Observed Events. However, when the intermediate events are erased, there is no operational procedure that allows one to access or verify the existence of a single joint probability distribution over all events in a given run. Assuming the existence of such a joint distribution therefore goes beyond what is operationally warranted. This motivates the assumption of Existence of Marginals.
At the same time, it remains operationally natural to require that, once the relevant pseudo events in the causal past are specified, future observed outcomes are screened off from additional variables. In this sense, pseudo events can be viewed as revealing intrinsic properties of the system that constrain what future events can or cannot occur. This requirement is captured by Operational Mediation and expresses a minimal notion of causal screening that does not rely on underlying ontic states or the full causal Markov condition. The following toy example explicitly demonstrates the legitimacy of a situation where Existence of Marginals and Operational Mediation can hold while Absoluteness of Observed Events fails.
Example: Existence of Marginals and Operational Marginals without Absoluteness of Observed Events. Consider the pseudo events $c,d\in\{0,1\}$, and Truly-Observed Events $a,b\in\{0,1\}$, and inputs $x,y\in\{0,1\}$. Let the pseudo events be distributed as $p(c,d)=1/4$, independent of $x,y$. Alice’s response functions are defined by $p(a=c|x=0,c)=1$ which depends on $c$ and $p(a=0|x=1,c)=p(a=1|x=1,c)=1/2$, independent of $c$. Bob’s response functions are defined analogously by $d$-dependent probability $p(b=d|y=0,d)=1$ and $p(b=0|y=1,d)=p(b=1|y=1,d)=1/2$, independent of $d$.
All operationally relevant marginals $p(c,d)$, $p(a|x,c)$, and $p(b|y,d)$ are well defined, and the pseudo events screen off future outcomes, so that $p(b|a,c,d,x,y)=p(b|y,d)$ and $p(a|b,c,d,x,y)=p(a|x,c)$, satisfying EOM and OM. However, because the outcomes associated with $x=0$ and $x=1$ (and similarly $y=0$ and $y=1$) cannot be jointly accessed in a single run due to the possibility of unitary erasure of pseudo events, the operational data do not warrant the existence of a single joint distribution $p(c,a,d,b|x,y)$. For instance, the relevant marginals can come from a joint quasi-probability distribution [64, 32]. Thus in this toy example Absoluteness of Observed Events is not satisfied, showing that Operational Pseudo Event Mediation is strictly weaker than it.
@@
It is important to note that the assumption Operational mediation is stronger than the Screening via Pseudo Events assumption used in the proof of the Causal Time-Symmetric Friendliness no-go theorem. In other words, Screening via Pseudo Events can be viewed as a specific instance of Operational mediation, where the dependence of the event $b$ on the past input choice $x$ is not relaxed. In the proof of Causal Time-Symmetric Friendliness no-go theorem, however, the independence of $b$ from $x$ is derived from the joint assumptions of Absoluteness of Observed Events, Axiological Time Symmetry, and No Retrocausality. In contrast, in the present analysis, we are not permitted to assume Absoluteness of Observed Events at the outset, and consequently, the independence of $b$ from $x$ is incorporated directly into the relatively stronger assumption of Operational mediation itself.
@@
Importantly, although Existence of marginals by itself does not require the existing marginals to originate from a single global joint probability distribution, imposing the requirement that those existing marginals given by Existence of marginals originate from a single global joint probability distribution leads to Absoluteness of Observed Events. Then we clearly have,
Absoluteness of Observed Events
$\implies$ Existence of marginals (51)
However, the reverse implication may not hold in general as Existence of marginals does not necessarily involve a single factorizable joint probability distribution.